### v2

# Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective

Abstract
We study the Yang-Lee theory in quantum phase transitions from the perspective of quantum entanglement in one-dimensional many-body systems. We primarily focus on the distribution of Yang-Lee zeros and its associated Yang-Lee edge singularity of two prototypical models: the Su-Schrieffer-Heeger model and the XXZ spin chain. By taking the zero-temperature limit, we show how the Yang-Lee zeros approach the quantum phase transition points on the complex plane of parameters. To characterize the edge singularity induced by Yang-Lee zeros in quantum phase transition, we introduce the entanglement entropy of the ground state to show that the edges of Yang-Lee zeros lead to the ground-state entanglement transition. We further show that our results are also applicable to the general non-interacting parity-time-symmetric Hamiltonians.
I Introduction
Yang-Lee zeros [1, 2], defined as the zero points of the partition function of the canonical ensemble, give a mathematical explanation of the nonanalyticity of thermodynamic observables at phase transition points. In the classical ferromagnetic Ising model, Yang and Lee investigated the zeros of the partition function in the presence of an imaginary magnetic field in order to understand the origin of the singularity in the ferromagnetic phase transition when one increases the temperature. The thermal phase transition takes place when Yang-Lee zeros touch the positive real axis on the complex plane of fugacity in the thermodynamic limit. In the vicinity of the edge of the distribution of Yang-Lee zeros, critical phenomena collectively known as Yang-Lee edge singularity emerge, which are accompanied by anomalous scaling laws [3, 4, 5, 6, 7, 8, 9, 10, 11]. Yang-Lee theory is applicable to a variety of thermal phase transitions in classical [12, 13, 14, 15] and quantum [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] systems.
Despite the great success of the application of Yang-Lee zeros to explain the origin of the singularity in thermal phase transitions, there exist exceptions in quantum phase transitions such as the topological quantum phase transition [29, 30], which are characterized by the change of the topological invariants instead of singularity in thermodynamic quantities. With the introduction of quantum entanglement, it was discovered that the singularity in a quantum phase transition can be more universally characterized by the singular behavior of the entanglement pattern in ground-state wavefunctions [31, 32, 33, 34]. Therefore, it is expected to universally characterize the Yang-Lee edge singularity in terms of the singularity in quantum entanglement instead of thermodynamic observables in the traditional Yang-Lee theory.
In this work we develop a Yang-Lee theory for quantum phase transitions and focus on how the Yang-Lee zeros approach the phase transition points by taking the zero-temperature limit. First, we investigate the distribution of Yang-Lee zeros in the parameter space of a noninteracting model: the Su-Schrieffer-Heeger (SSH) model [35] by generalizing the Rice-Mele coupling [36] to an imaginary one. We explore the nonunitary criticality of the non-Hermitian SSH model originating from Yang-Lee edge singularity. These quantum critical phenomena are induced by the exceptional points where a nonanalytic excitation spectrum exists. Furthermore, we figure out that the edges of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground states. We also generalize our discussion to general parity-time (PT)-symmetric free-fermion systems to unveil universally unnoticed universality in Yang-Lee edge singularity in terms of transitions in the entanglement entropy. This result indicates that entanglement entropy may serve as a good characterization for the universality class of Yang-Lee edge singularity in quantum phase transitions. In contrast to Ref. [34], we have rigorously proved the correspondence between the edges of Yang-Lee zeros and ground-state entanglement transitions for all PT-symmetric free-fermion systems rather than entanglement transitions of the long-time steady states of a class of specific models.
We also discuss the Yang-Lee theory of an interacting model: the XXZ spin chain. On the complex plane of the anisotropy, we investigate the distribution of Yang-Lee zeros near the ferromagnetic phase transition point. By analyzing the Bethe ansatz of the XXZ model, we prove that the Yang-Lee zeros approach the phase transition point in a line for arbitrary system size. The edge of the distribution of Yang-Lee zeros turns out to be the ferromagnetic phase transition point. We also show that the edge of the Yang-Lee zeros corresponds to an entanglement transition. This exact result paves the way for figuring out Yang-Lee zeros in quantum phase transitions of interacting models. Finally, we provide a general argument for the correspondence between Yang-Lee edge singularity and entanglement transition.
II Non-interacting Model
We begin from a non-interacting model: the non-Hermitian SSH model. The Hamiltonian in the momentum space is defined as [37, 35]
$H=\sum_{k}(c_{kA}^{\dagger}\ \ c_{kB}^{\dagger})H_{k}\left(\begin{array}[]{cc}c_{kA}\\ c_{kB}\end{array}\right),$ (1)
with the Bloch Hamiltonian
$H_{k}=\left(\begin{array}[]{cc}iu&we^{-ik}+v\\ we^{ik}+v&-iu\end{array}\right),$ (2)
where $k$ is the momentum and $c_{kA}^{(\dagger)}$ and $c_{kB}^{(\dagger)}$ represent the annihilation (creation) operators of the fermions with momentum $k$ on the sublattice $A$ and $B$ respectively. The energy spectrum of the Bloch Hamiltonian is $E_{k}^{\pm}=\pm\sqrt{|v+we^{ik}|^{2}-u^{2}}=:\pm E_{k}$. The model has three phases: the trivial PT-unbroken gapped phase for $w-v<-u$, the PT-broken gapless phase for $-u<w-v<u$, and the topologically nontrivial PT-unbroken gapped phase for $w-v>u$ [35].
Figure 1: Distribution of Yang-Lee zeros (shaded region) of the SSH model on the plane of the temperature and the hopping parameters. At zero temperature, the onset of Yang-Lee zeros coincides with the quantum phase transitions indicated by the band structure (blue curves).
The partition function of the non-Hermitian SSH model is defined as  [^1] Here we assume the chemical potential $\mu=0$.
$Z$ $:=\sum_{n}\langle E_{n}^{L}|e^{-\beta H}|E_{n}^{R}\rangle.$
$=\prod_{k}(1+e^{-\beta E_{k}})(1+e^{\beta E_{k}}),$ (3)
where $\beta:=1/k_{B}T$ is the inverse temperature and the states $|E_{n}^{L(R)}\rangle$ are the left (right) eigenstates of the Hamiltonian with the energy $E_{n}$. The conditions of zeros of the partition function can be determined as
$\text{Re}[E_{k}]=0,\text{Im}[E_{k}]=(2n+1)\pi/\beta$ (4)
with $n\in\mathbb{Z}$. To figure out the distribution of Yang-Lee zeros, we first assume a finite $\beta$ and investigate where the condition (4) is satisfied in the momentum space when we take the thermodynamic limit. Then we take the zero-temperature limit to see how the distribution of Yang-Lee zeros approaches the quantum phase transition points. The order of taking the two limits cannot be exchanged. In the zero-temperature limit $\beta\to\infty$, we find that the condition (4) is satisfied in the thermodynamic limit for the gapless phase $|v-w|\leq u$, corresponding to the whole PT-broken phase. In the gapless phase, there always exists an interval in the momentum space with a finite purely imaginary part of the energy spectrum, which induces the Yang-Lee zeros. In Fig. 1, we show how Yang-Lee zeros approach the quantum phase transition points $|w-v|=u$ by taking the zero-temperature limit. To further understand the distribution of Yang-Lee zeros, we introduce the number of roots of the partition function $\chi$ [26], which is defined as the number of the integer $n$ satisfying the condition of zeros in Eq. (4). The number of roots $\chi$ can be understood as the density of Yang-Lee zeros on the plane of parameters. In the zero-temperature limit, the number of roots $\chi$ takes the form of
$\chi/\beta\simeq\frac{1}{2\pi}\sqrt{u^{2}-|v-w|^{2}},$ (5)
where $u>|v-w|$ in the PT-broken phase. Near the phase boundary $|v-w|=u$, the asymptotic behavior of $\chi$ can be shown as $\chi\propto(u-|v-w|)^{1/2}$ which is attributed to the exceptional points [26] in the gapless phase. Since the dispersion relation is $E_{k}^{\pm}=\pm\sqrt{v^{2}+w^{2}+2vw\cos k-u^{2}}$, the gap closes at
$k_{E}=\pm\arccos\frac{u^{2}-v^{2}-w^{2}}{2vw},$ (6)
which correspond to the exceptional points. Around the exceptional points, the excitation spectrum takes the form of $E\propto(k-k_{E})^{1/2}$, giving rise to the asymptotic behavior of $\chi$.
Furthermore, we examine the Yang-Lee edge singularity by the correlation functions and its associated critical exponents. The correlation function
$C_{\alpha\beta}(x)$ $={}_{L}\langle c_{\alpha}^{\dagger}(x)c_{\beta}(0)\rangle_{R},$
$:=\frac{1}{Z}\langle G_{L}|c_{\alpha}^{\dagger}(x)c_{\beta}(0)e^{-\beta H}|G_{R}\rangle,$ (7)
where $\alpha,\beta\in\{A,B\}$ indicating the sublattice index and the states $|G_{L(R)}\rangle$ represents the left (right) eigenstates of the Hamiltonian (1) with only the negative real-energy modes filled. In the gapped phases $|v-w|>u$, this correlation function can be exactly calculated near the phase boundary at zero temperature as
$\lim_{x\to\infty}C_{\alpha\beta}(x)\propto\frac{e^{-x/\xi}}{\sqrt{x}},$ (8)
where $\delta=|v-w|-u$ and $\xi=vw/(2u\delta)$ is the correlation length. See Appendix A for details. Hence, from Eq. (8), we obtain the critical exponent $\nu=1$ which is defined as $\xi^{-1}\propto\delta^{\nu}$ [38]. In addition, from the correlation function (8), we have the anomalous scaling dimension $\eta=3/2$ defined as $\lim_{x\to\infty}C(x)\propto\exp(-x/\xi)/x^{d-2+\eta}$ [38]. The anomalous power $1/2$ originates from the dispersion relation $E\propto(k-k_{E})^{1/2}$ near the exceptional points. Hence, combining with the dynamical critical exponent $z=1$, we have the critical exponents for the Yang-Lee universality class here, which are similar to the ones in the Yang-Lee universality class in the non-Hermitian BCS superconductivity [26]. They are both induced by exceptional points in a PT-symmetric Hamiltonian.
Now we turn to discuss the singularity induced by Yang-Lee zeros in terms of the entanglement entropy (EE) of the ground state. The EE of the subsystem $A$ is defined as $S_{A}:=\text{Tr}_{A}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}=\text{Tr}_{\bar{A}}|G_{R}\rangle\langle G_{L}|$ [35]. The EE discontinuously jumps from a subarea-law scaling $S_{A}=(1/6)\ln L_{A}$ in the PT-broken gapless phase, where Yang-Lee zeros are located, to an area-law scaling $S=\text{const}$ in the PT-unbroken gapped phase [39, 35]. Here $L_{A}$ denotes the length of the subsystem $A$. We can see that the edges of the distribution of zeros correspond to the entanglement transitions. We note that the correspondence is also applicable to the Hermitian SSH model. In the Hermitian case $u=0$, the system at the phase transition point $v=w$, which is the edge of Yang-Lee zeros, exhibits a subarea-law scaling for the ground-state EE. To elucidate the generality of the correspondence, we consider an arbitrary local PT-symmetric free-fermion Hamiltonian. By generalizing the discussion in Ref. [40], we prove that the gapless PT-broken phase with Yang-Lee zeros shows at least a subarea-law scaling $S_{A}\geq c_{-}\ln L_{A}$ with $c_{-}>0$ and the gapped PT-unbroken phase shows an area-law scaling in the one-dimensional case. Therefore, the correspondence between the edges of the distribution of Yang-Lee zeros and the entanglement transitions is general for PT-symmetric free-fermion Hamiltonians. The Yang-Lee zeros introduce the singularity to the entanglement transition. Thus, entanglement entropy can be considered as a universal characterization of the Yang-Lee edge singularity in quantum phase transitions. See Appendix. B for the details of the proof.
III Interacting Model
Above we have considered the Yang-Lee zeros in the quantum phase transitions of the non-interacting cases. In this part, we extend our discussion to an interacting case. We investigate the XXZ model to understand the emergence of Yang-Lee zeros from the Bethe ansatz. The Hamiltonian of the XXZ spin chain is defined as
$H=-J\sum_{i=1}^{L}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}+\Delta S_{i}^{z}S_{i+1}^{z}),$ (9)
where $J>0$, $L$ is the size of the spin chain and we extend the anisotropy parameter $\Delta$ to a complex one to consider the distribution of Yang-Lee zeros. Here we focus on the ferromagnetic phase transition point $\Delta=1$. In the Hermitian case, the ground state of the gapped (gapless) phase $\Delta>1$ ($-1<\Delta<1$) is (anti)ferromagnetic [41].
To consider the condition of Yang-Lee zeros, we follow the definition of the partition function in Eq. (II): $Z=\sum_{n}{}_{L}\langle E_{n}|\exp(-\beta H)|E_{n}\rangle_{R}$. We show the following theorem:
Theorem 1.
The solutions $\Delta$ of $Z=0$ around the ferromagnetic phase transition point $\Delta=1$ are given by the algebraic equation
$\sum_{M=0}^{L}z^{M(L-M)}=0$ (10)
in the zero-temperature limit $\beta\to\infty$ for an arbitrary length $L$ of the spin chain with $z:=\exp(\beta J(\Delta-1)/(L-1))$.
From Eq. (10), we deduce that overall, there are $N$ complex roots of $z$ which can be defined as $z_{1},\cdots,z_{N}$, where $N=(L/2)^{2}$ for even $L$ and $N=(L^{2}-1)/4$ for odd $L$. By rewriting Eq. (10) as $\sum_{M=1}^{N}a_{M}z^{M}=0$, we have $a_{0}=2$ and $a_{N}=1$ for an even $L$ or $a_{0}=a_{N}=2$ for an odd $L$, indicating that both $z=0$ and $z=\infty$ are not the solutions of Eq. (10). Therefore, all the roots $z$ are finite that give rise to the zeros $\Delta_{j}$ on the complex plane of $\Delta$ as
$\Delta_{j}=1+\frac{L-1}{\beta J}\log(z_{j})+i(L-1)\frac{2n\pi}{\beta J},$ (11)
where $n\in\mathbb{Z}$ and $j=1,\cdots,N$. We can see that the zeros form a line around the phase transition point. In the zero-temperature limit, those zeros are equivalent to $\Delta_{j}=1+i(L-1)\frac{2n\pi}{\beta J}$, which are simply the line $\text{Re}\Delta=1$. This indicates that the Yang-Lee zeros approach the real axis of $\Delta$ through the line $\text{Re}\Delta=1$. Below we outline the proof of Theorem 1.
Proof.
To begin with, we explain the structure of the exact solution of the XXZ spin chain. The Bethe ansatz equations (BAEs) of the XXZ chain are given by [42, 43, 41]
$\left(\frac{\sin\frac{\phi}{2}(x_{j}+i)}{\sin\frac{\phi}{2}(x_{j}-i)}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l}+2i)}{\sin\frac{\phi}{2}(x_{j}-x_{l}-2i)},$ (12)
where we define $x_{j}$ as $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$ and $\phi:=\cosh^{-1}\Delta$ with $k_{j}$ being the quasimomenta of the excited magnons. In Eq. (12), $j=1,\cdots,M$ where $M$ represents the number of magnons. The energy of the $M$-magnon sector takes the form of [41]
$E_{M}$ $=E_{0}+\sum_{j=1}^{M}J(\Delta-\cos k_{j})$
$=E_{0}+\sum_{j=1}^{M}\frac{J\sinh^{2}\phi}{\cosh\phi-\cos\phi x_{j}},$ (13)
where $E_{0}=-JL\Delta/4$ represents the ground state energy for the states with all spins up or down. Near the ferromagnetic phase transition point $\Delta=1$, we introduce a quantity $\delta:=\Delta-1$ and the quantity $\phi$ can be approximated as $\phi\approx\sqrt{2\delta}$. To understand the behavior of Yang-Lee zeros, we first turn to a limit $\delta\to 0$, under which the system becomes the Heisenberg model. We can see the $(L+1)$-degeneracy of the ground state in the Heisenberg model. Those ground states can be constructed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where we denote $|\Omega\rangle$ as the all-spin-up state. However, the $(L+1)$-degeneracy will be lifted if we add a perturbation to the parameter $\Delta$. To analyze the gap between levels, we consider the expansion of the BAEs for XXZ chain (12) around $\phi=0$,
$L\cot\frac{\phi}{2}x_{j}=2\sum_{l\neq j}\cot\frac{\phi}{2}(x_{j}-x_{l}),$ (14)
which we call simplified Bethe ansatz equations (SBAEs) and are shown to be a set of algebraic equations of $\cot(\phi x_{j}/2)$. From the equations, one can prove that the solutions $\cot(\phi x_{j}/2)$ only depend on the size $L$. Meanwhile, the energy can also be expressed as a function of $\cot(\phi x_{j}/2)$,
$E_{M}=E_{0}+\sum_{j=1}^{M}\frac{2J\delta(1+\cot^{2}(\phi x_{j}))}{2+\delta\cot^{2}(\phi x_{j}/2)}.$ (15)
By substituting the solutions of the SBAEs into Eq. (15), we obtain the energy of the $M$-magnon sector as
$E_{M}=E_{0}+J\delta\frac{M(L-M)}{L-1}.$ (16)
With the definition $z=\exp(\beta J\delta/(L-1))$, we can rewrite the condition of $Z=0$ with Eq. (16) as
$\sum_{M=0}^{L}z^{M(L-M)}=0,$ (17)
which is equivalent to Eq. (10). The maximal value of the exponent $M(L-M)$ is $L^{2}/4$ for an even $L$ with $M=L/2$ or $(L^{2}-1)/4$ for an odd $L$ with $M=(L-1)/2,(L+1)/2$. This completes the proof. Details of the proof are given in Appendix. C
To numerically verify the line distribution of the Yang-Lee zeros, we show the zeros of the partition function on the complex plane in Fig. 2(a) for $L=6$.
Figure 2: (a): The distribution line of Yang-Lee zeros in the non-Hermitian XXZ model on the complex plane of $\Delta$. The size of the spin chain is $L=6$ and the inverse temperature is $\beta=100$. The zeros of the partition function approach the phase transition point $\Delta=1$ through a straight line. The dashed lines represent the entanglement between spins. Here we show the nearest 200 zeros to the phase transition point. (b): The scaling of EE $S_{A}$ of subsystem $A$ with size $L_{A}$ at the point $\Delta=0.99+0.01i$ where $L=200$. The fitting curve is chosen as $S_{A}=1.881+0.184\log(\sin(\pi L_{A}/L)))$, exhibiting the subarea law.
We can further understand the line distribution of Yang-Lee zeros from the gap behavior of the non-Hermitian XXZ model on the complex plane. For the Hamiltonians with $\text{Re}\Delta>1$, we can see that the systems must be gapped from Eq. (16), with the gap $\Delta E=J\text{Re}\delta$. However, for those Hamiltonians with $\text{Re}\Delta<1$, the systems become gapless. In this phase, the ground state has $M=L/2$ magnons and the real part of the energy becomes $\text{Re}E=E_{0}+\frac{(L/2)^{2}}{L-1}J\text{Re}\delta$ with $\text{Re}\delta<0$ when $L$ is even. Then the energy level distance between the ground state and the first excited state is given by
$\Delta E=-\frac{J\text{Re}\delta}{L-1},$ (18)
which vanishes in the thermodynamic limit. Hence, the Hamiltonians with $\text{Re}\Delta<1$ behave as gapless systems. The behavior of the gap supports the existence of the Yang-Lee zeros near the phase transition point $\Delta=1$.
Next, we examine the critical exponents for the Yang-Lee edge singularity of the non-Hermitian XXZ model. We introduce a magnetic field $h$ to the Hamiltonian. By solving the BAE (12), the susceptibility in the gapless phase takes the form of $\lim_{h\to 0}\chi=\frac{2\gamma}{J\pi\delta}$ near the critical point where $\gamma$ is a constant [41]. Therefore, we obtain the critical exponent $\sigma=1$ from $\chi$ defined as $\chi\sim\delta^{-\sigma}$. At the critical point, the susceptibility diverges. In addition, by definition of the density of Yang-Lee zeros $g(\Delta)=\sum_{j}\delta(\Delta-\Delta_{j})$, we have
$g=\beta J\frac{N}{2\pi(L-1)}.$ (19)
at the phase transition point. In the thermodynamic limit, the density has the scaling $g=\beta JL/8\pi$.
Furthermore, here the edge of the Yang-Lee zeros is also associated with the entanglement transition of the ground state. The zeros are distributed on the line $\text{Re}\Delta=1$ and approach the phase transition point $\Delta=1$ when we decrease the temperature. Hence, we can consider the phase transition point as the edge of the Yang-Lee zeros. On the real axis, the EE of the gapless region $\text{Re}\Delta<1$ shows a subarea-law scaling while that of the gapped region $\text{Re}\Delta>1$ shows an area-law scaling. Therefore, the correspondence between the Yang-Lee edge singularity and the entanglement transitions is applicable to not only the non-interacting cases but also the interacting systems. We note that the points on the distribution line of the Yang-Lee zeros also represent the entanglement transition points. To illustrate this, we calculate the EE of the model at $\Delta=0.99+0.01i$ with the non-Hermitian density-matrix renormalization group [44, 45] in Fig. 2(b), which also shows a subarea-law scaling. The reduced density matrix of the subsystem $A$ is defined as $\rho_{A}=\text{Tr}_{\bar{A}}|\psi_{0}^{R}\rangle\langle\psi_{0}^{R}|$, where $|\psi_{0}^{R}\rangle$ is the right eigenstate of the Hamiltonian (9) with the lowest real energy. On the other hand, the systems with $\text{Re}\Delta>1$ show the area law in the gapped phase since the ground state in this phase is a direct-product state. Therefore, all of the Yang-Lee zeros correspond to entanglement transition points on the complex plane.
To generally understand how Yang-Lee zeros introduce the singularity in the entanglement transitions, we consider the density matrix $\rho=e^{-\beta H}/Z$ for the Hamiltonian $H(\lambda)$ with $\lambda$ being a parameter and assume the critical point of the system to be $\lambda=\lambda_{c}$ where we have the entanglement transition. We generalize the parameter $\lambda$ into a complex one and consider the Yang-Lee zeros on the complex plane of $\lambda$. Let the EE be $S_{A}=-\text{Tr}_{A}[\rho_{A}\log\rho_{A}]$ where $\rho_{A}=\text{Tr}_{\bar{A}}\rho$ is the reduced density matrix of the subsystem $A$ at absolute zero. Since the zeros of the partition function also represent singularity of the EE, once one can find Yang-Lee zeros in the vicinity of $\lambda=\lambda_{c}$ on the complex plane of $\lambda$ at absolute zero, the singularity can enter the Hermitian system at $\lambda=\lambda_{c}$ and hence the entanglement transition takes place. The critical point $\lambda=\lambda_{c}$ is the edge of the distribution of Yang-Lee zeros.
IV Conclusion and Outlook
In this work, we have investigated the distribution of Yang-Lee zeros in quantum phase transitions of the SSH model and the XXZ model and shown how Yang-Lee zeros enter the phase transition points. In particular, in the XXZ model, we have proposed the simplified Bethe ansatz equation (SBAE) to exactly solve the energy levels around the ground state and calculated the partition function in the zero-temperature limit. We have also explored the critical behavior at the edges of Yang-Lee zeros of both models. Furthermore, we find that the edges of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground state and discuss the mechanism behind the correspondence. We have also generalized the correspondence to an arbitrary PT-symmetric free-fermionic system.
The Yang-Lee zeros in quantum phase transitions studied in this paper are not only an interesting mathematical concept but also experimentally realizable. The non-Hermitian SSH model can be realized by dilating it into a larger Hermitian Hamiltonian due to the PT symmetry [46, 47, 48]. We can introduce an ancillary qubit as an environment coupled to the targeted spin to realize the Hamiltonian (2). For the non-Hermitian XXZ model, we can realize it in a non-Hermitian Bose-Hubbard model in a strongly correlated regime [44, 49, 50], where the non-Hermiticity originates from the two-body loss in the open quantum systems. By coupling the target system to an ancilla, we expect the Yang-Lee zeros and its associated edge singularity to be measurable with the decoherence of the spin in the ancilla [51, 52, 53, 19]. The Yang-Lee zeros in those systems can also be measured from the diagonal elements of the time-evolution operator $U=\exp(i\tilde{H}t)$, where $\tilde{H}=iH$ and $t=\beta$ [54]. To be specific, we can measure the partition function of a non-Hermitian SSH model by preparing a state whose dynamics is described by the $2\times 2$ Bloch Hamiltonian $iH_{k}$ in the momentum space, where $H_{k}$ is given in Eq. (2). Then the trace of the evolution operator gives the partition function: $Z=\prod_{k}\text{Tr}U_{k}$ with $t=\beta$ and $U_{k}=\exp{[i(iH_{k})t]}=\exp{(-H_{k}t)}$. Similarly, we can also prepare a state whose dynamics is governed by the Hamiltonian $iH$ with tomography [55] to measure the partition function of the non-Hermitian XXZ model where $H$ is shown in Eq. (9). The partition function is given by $Z=\text{Tr}U=\text{Tr}\exp{[i(iH)t]}=\text{Tr}\exp{(-Ht)}$ where $t=\beta$. The transition of the entanglement entropy is expected to be observed in the quantum gas microscope [56].
V ACKNOWLEDGMENTS
We are grateful to Zongping Gong, Masaya Nakagawa, Shinsei Ryu, Yunfeng Jiang, Kohei Kawabata, and Masahito Ueda for fruitful discussion. H. L. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, at the University of Tokyo. This work was also supported by JSPS KAKENHI Grants No. 24KJ0824.
Appendix A Correlation Function in the SSH model
Here we consider the Yang-Lee zeros in the non-Hermitian SSH model. The Bloch Hamiltonian of non-Hermitian SSH model is given by [35]
$H_{k}=iu\sigma_{z}+(v+w\cos k)\sigma_{x}+w\sin k\sigma_{y}=\left(\begin{array}[]{cc}iu&we^{-ik}+v\\ we^{ik}+v&-iu\end{array}\right).$ (20)
The energy spectrum is $E_{k,\pm}=\pm\sqrt{|we^{-ik}+v|^{2}-u^{2}}\equiv\pm E_{k}$. Overall, there are two phases in the Hamiltonian: the PT-symmetric phase for $|w-v|>u$ and the PT-broken phase for $|w-v|<u$. In the PT-symmetric phase, the Hamiltonian is gapped with the energy gap $\Delta=2\sqrt{(w-v)^{2}-u^{2}}$. In the PT-broken phase, the Hamiltonian is gapless with $\text{Re}[E_{k}]=0$ between the exceptional points given by
$k_{E}=\pm\arccos\frac{u^{2}-v^{2}-w^{2}}{2vw},$ (21)
where the gap closes and the Hamiltonian cannot be diagonalized. The excitation spectrum around the exceptional points is given by $E_{k}\sim\sqrt{|k-k_{E}|}$ [35].
Now we turn to calculate the critical exponents around the phase transition points. We consider the two-point correlation function. To begin with, we first diagonalize the Hamiltonian as
$H$ $=$ $\sum_{k}\left(\begin{array}[]{cc}c_{kA}^{\dagger}&c^{\dagger}_{kB}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{c}c_{kA}\\ c_{kB}\end{array}\right)$ (27)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&\sin\frac{\phi_{k}}{2}\\ -\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\\ \sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right)$ (37)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}E_{k}&0\\ 0&-E_{k}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right),$ (43)
where the transformation is given by
$\gamma_{kA}$ $=$ $\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}+\sin\frac{\phi_{k}}{2}c_{kB},$ (44)
$\gamma_{kB}$ $=$ $-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB},$ (45)
$\bar{\gamma}_{kA}$ $=$ $\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},$ (46)
$\bar{\gamma}_{kB}$ $=$ $-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB}.$ (47)
From the transformation, we have $\bar{\gamma}_{kA}\neq\gamma_{kA}^{\dagger}$ and $\bar{\gamma}_{kB}\neq\gamma_{kB}^{\dagger}$ since the Hamiltonian is non-Hermitian. However, the anti-commutation relations between $\gamma_{kA(B)}$ and $\bar{\gamma}_{kA(B)}$ still follow the fermionic ones, which can be proved as below:
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}A}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (48)
$=$ $\cos^{2}\frac{\phi_{k}}{2}\{a_{kA}^{\dagger},a_{k^{\prime}A}\}+\sin^{2}\frac{\phi_{k}}{2}\{a_{kB}^{\dagger},a_{k^{\prime}B}\}$
$=$ $\delta_{k,k^{\prime}},$
$\{\bar{\gamma}_{kB},\gamma_{k^{\prime}B}\}$ $=$ $\left\{-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (49)
$=$ $\sin^{2}\frac{\phi_{k}}{2}\{c_{kA}^{\dagger},c_{k^{\prime}A}\}+\cos^{2}\frac{\phi_{k}}{2}\{c_{kB}^{\dagger},c_{k^{\prime}B}\}$
$=$ $\delta_{k,k^{\prime}},$
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}B}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB}\right\}$ (50)
$=$ $-\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\{c_{kA}^{\dagger},c_{k^{\prime}A}\}+\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\{c_{kB}^{\dagger},c_{k^{\prime}B}\}$
$=$ $0$
Hence, we can still consider $\bar{\gamma}_{kA(B)}$ as the creation operator of the fermion with the flavor $A(B)$ and the momentum $k$, similar to the non-Hermitian Bogoliubov transformation in Ref. [57]. Meanwhile, we can calculate the correlation functions with the linear transformations (44)-(47). Here we define the correlation function as
$C_{\alpha\beta}(x)=_{L}\langle c_{\alpha}^{\dagger}(x)c_{\beta}(0)\rangle_{R}:=\frac{1}{Z}\langle G_{L}|c_{\alpha}^{\dagger}(x)c_{\beta}(0)e^{-\beta E_{ks}}|G_{R}\rangle,$ (51)
where $\alpha,\beta\in\{A,B\}$ and the state $|G_{L(R)}\rangle$ represent the right (left) eigenstate of the Hamiltonian with only the negative real-energy modes filled. It has four components which can be calculated from the Fourier transformation. We first show the forms of correlation functions in the momentum space as
$C_{AA}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kA}\rangle_{R}$ (52)
$=$ (${}_{L}\left\langle\left(\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\bar{\gamma}_{kA}-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\bar{\gamma}_{kB}\right)\left(\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\gamma_{kA}-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\gamma_{kB}\right)\right\rangle_{R}$
$=$ $\cos^{2}\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kA}\gamma_{kA}\rangle_{R}+\sin^{2}\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kB}\gamma_{kB}\rangle_{R}$
$=$ $\cos^{2}\frac{\phi_{k}}{2}\frac{1}{e^{\beta E_{k}}+1}+\sin^{2}\frac{\phi_{k}}{2}\frac{1}{e^{-\beta E_{k}}+1},$
$C_{BB}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kB}\rangle_{R}$ (53)
$=$ (${}_{L}\left\langle\left(\sin\frac{\phi_{k}}{2}\bar{\gamma}_{kA}+\cos\frac{\phi_{k}}{2}\bar{\gamma}_{kB}\right)\left(\sin\frac{\phi_{k}}{2}\gamma_{kA}+\cos\frac{\phi_{k}}{2}\gamma_{kB}\right)\right\rangle_{R}$
$=$ $\sin^{2}\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kA}\gamma_{kA}\rangle_{R}+\cos^{2}\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kB}\gamma_{kB}\rangle_{R}$
$=$ $\sin^{2}\frac{\phi_{k}}{2}\frac{1}{e^{\beta E_{k}}+1}+\cos^{2}\frac{\phi_{k}}{2}\frac{1}{e^{-\beta E_{k}}+1},$
$C_{AB}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kB}\rangle_{R}$ (54)
$=$ (${}_{L}\left\langle\left(\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\bar{\gamma}_{kA}-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\bar{\gamma}_{kB}\right)\left(\sin\frac{\phi_{k}}{2}\gamma_{kA}+\cos\frac{\phi_{k}}{2}\gamma_{kB}\right)\right\rangle_{R}$
$=$ $\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kA}\gamma_{kA}\rangle_{R}-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\cos\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kB}\gamma_{kB}\rangle_{R}$
$=$ $\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\left(\frac{1}{e^{\beta E_{k}}+1}-\frac{1}{e^{-\beta E_{k}}+1}\right),$
$C_{BA}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kA}\rangle_{R}$ (55)
$=$ (${}_{L}\left\langle\left(\sin\frac{\phi_{k}}{2}\bar{\gamma}_{kA}+\cos\frac{\phi_{k}}{2}\bar{\gamma}_{kB}\right)\left(\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\gamma_{kA}-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\gamma_{kB}\right)\right\rangle_{R}$
$=$ $\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kA}\gamma_{kA}\rangle_{R}-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\cos\frac{\phi_{k}}{2}{}_{L}\langle\bar{\gamma}_{kB}\gamma_{kB}\rangle_{R}$
$=$ $\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\left(\frac{1}{e^{\beta E_{k}}+1}-\frac{1}{e^{-\beta E_{k}}+1}\right),$
Here we only consider the LR correlation function and set the chemical potential $\mu=0$. Then we consider the correlation function in the real space by Fourier transformation. In the gapped phase we have
$C_{AA}(x)$ $=$ $\int\frac{dk}{2\pi}\langle c_{kA}^{\dagger}c_{kA}\rangle e^{ikx}$ (56)
$=$ $\int\frac{dk}{2\pi}\left(\cos^{2}\frac{\phi_{k}}{2}\frac{1}{e^{\beta E_{k}}+1}+\sin^{2}\frac{\phi_{k}}{2}\frac{1}{e^{-\beta E_{k}}+1}\right)e^{ikx}$
$\xrightarrow{T=0}$ $\int\frac{dk}{2\pi}\frac{1}{2}\left(1-\frac{iu}{\sqrt{|v_{k}|^{2}-u^{2}}}\right)e^{ikx}$
[…]
This integral can approximately be calculated near the critical point. Near the critical points $|v-w|=u$, we have
[…]
In this integral, we find that the parts around the points $k=\pm\pi$ dominate. Hence, we expand around the two points and obtain  [^2] Here we only take the critical point $v-w=u$ as an example.
[…]
$=$ $-e^{i\pi x}\frac{iu}{2\pi\sqrt{2u\delta}\xi}K_{0}\left(\frac{x}{\xi}\right),$
where $\delta=v-w-u$ and $\xi=vw/2u\delta$. Here $K_{0}(x)$ denotes the modified Bessel function of the second kind. Similarly, we also have
[…]
$=$ $e^{i\pi x}\frac{iu}{2\pi\sqrt{2u\delta}\xi}K_{0}\left(\frac{x}{\xi}\right).$
$C_{AB}(x)$ $=$ $\int\frac{dk}{2\pi}\left[\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\left(\frac{1}{e^{\beta E_{k}}+1}-\frac{1}{e^{-\beta E_{k}}+1}\right)\right]e^{ikx}$ (60)
$\xrightarrow{T=0}$ $-\frac{1}{2}\int\frac{dk}{2\pi}\sin\phi_{k}e^{ikx}$
$=$ $-\frac{1}{2}\int\frac{dk}{2\pi}\sqrt{1+\frac{u^{2}}{v^{2}+w^{2}+2vw\cos k-u^{2}}}e^{ikx}$
$\simeq$ $-\int_{-\infty}^{\infty}\frac{dk}{2\pi}\sqrt{1+\frac{u/2\delta}{1+\xi^{2}k^{2}}}\cos(kx)$
$\simeq$ $-\frac{1}{2\pi\xi}\sqrt{\frac{2\delta}{u}}K_{0}\left(\frac{x}{\xi}\right).$
The correlation function $C_{BA}(x)$ shows the same result as $C_{AB}(x)$. Therefore, all the correlation functions have the same long-range scaling behavior with the distance $x$ as
$\lim_{x\rightarrow\infty}C_{\alpha\beta}(x)\sim\frac{e^{-x/\xi}}{\sqrt{x}},$ (61)
which originates from the scaling behavior of $K_{0}(x)$. This exponential decay is attributed to the finite energy gap of the system. From the definition of the critical exponents in the main text, we obtain $\nu=1,\eta=3/2$.
Appendix B Entanglement Entropy in Non-interacting Fermionic Systems
In this section, we calculate the entanglement entropy for the general local non-interacting fermionic systems in the $d$-dimensional case with PT symmetry. Here we consider the entanglement entropy of the ground-state density matrix $\rho=|\psi_{0}^{R}\rangle\langle\psi_{0}^{L}|$, where $|\psi_{0}^{R(L)}\rangle$ represents the right (left) ground state of the system. The entanglement entropy of the subsystem $A$ is defined as
$S_{A}=-\text{Tr}[\rho_{A}\log\rho_{A}],\rho_{A}:=\text{Tr}_{\bar{A}}\rho.$ (62)
The ground state here is defined as the eigenstate with only the negative-energy modes filled. Here we consider the case of ground-state degeneracy where the system has two phases: the PT-broken phase and the PT-unbroken phase. The system has pure imaginary eigenvalues in the PT-broken phase which indicates that the phase is gapless and only finite real values in the PT-unbroken phase which indicates that the phase is gapped. In the PT-broken phase, the pure imaginary eigenvalues are surrounded by the exceptional points in the Brillouin zone, namely exceptional rings [58, 59]. The partition function defined in Eq. (5) in the main text can be written as
$Z=\prod_{k,\alpha}\left(1+e^{-\beta E_{k}^{\alpha}}\right),$ (63)
where $\alpha$ is the index of band. In the PT-broken phase, since $E_{k}^{\alpha}$ can be imaginary values for some given $\alpha$, we have $Z=0$ for those momenta satisfying
$\text{Im}E_{k}^{\alpha}=\frac{(2n+1)\pi}{\beta}.$ (64)
We can infer that there always exist momenta $k$s satisfying Eq. (64) at absolute zero in the PT-unbroken phase, where Yang-Lee zeros emerge. Next, we are going to prove that the entanglement entropy shows at least a subarea-law scaling in the PT-unbroken phase. To show this, we follow the method used in Ref. [40]. The correlation matrix is defined as
$\gamma_{i\alpha,j\beta}=\delta_{ij}\delta_{\alpha\beta}-2\text{Tr}[\rho c^{\dagger}_{i\alpha}c_{j\beta}]=\langle\psi_{0}^{L}|(I_{ij}I_{\alpha\beta}-2c^{\dagger}_{i\alpha}c_{j\beta})|\psi_{0}^{R}\rangle,$ (65)
where $i,j\in\mathbb{Z}$ are the positions of the fermions in the subsystem $A$, $\alpha,\beta$ are the band indices and $I$ is the identity matrix. To simplify the calculation, here we assume that the subsystem is a hypercube with edge length $L$. The entanglement entropy can be expressed with the eigenvalues $\lambda_{M}$ of the correlation matrix in the subsystem $A$ [35].
$S_{A}$ $=\sum_{M=1}^{NL^{d}}h(\lambda_{M}),$ (66)
$h(x)$ $=-\frac{1+x}{2}\log\left(\frac{1+x}{2}\right)-\frac{1-x}{2}\log\left(\frac{1-x}{2}\right),$ (67)
where $N$ is the number of the energy bands. We begin from a two-band case. We consider the case where the Fermi surface is the exceptional ring and the correlation matrix is given by
$\gamma_{i\alpha,j\beta}=\int d^{d}k(1-2\eta(k))e^{ik(i-j)}\delta_{\alpha\beta},$ (68)
where $\eta(k):=\theta(-\text{Re}E_{k}+\epsilon)$ is an indicator function with $\epsilon>0$ an infinitesimal number. We note that here we choose the right and left ground states as $|\psi_{0}^{R}\rangle=\prod_{k:\text{Re}[E_{k}^{\alpha}]<0}c_{k\alpha}^{\dagger}|\mathrm{vac}\rangle$ and $|\psi_{0}^{L}\rangle=\prod_{k:\text{Re}[E_{k}^{\alpha}]<0}\langle\mathrm{vac}|c_{k\alpha}$. By utilizing the inequality
$h(x)>\frac{1}{2}(1-x^{2}),$ (69)
we have the lower bound of the entanglement entropy
$S_{A}>\frac{1}{2}\text{Tr}[I-\gamma^{2}].$ (70)
With the introduction of the kernel [40]
$F_{L}(x)=\sum_{i,j\in\mathbb{Z}_{L}}e^{ix(i-j)}=\frac{\cos(Lx)-1}{\cos(x)-1},$ (71)
we have
$\text{Tr}[I-\gamma^{2}]$ $=\frac{4}{(2\pi)^{2d}}\int d^{d}k\int d^{d}k^{\prime}\eta(k)(1-\eta(k^{\prime}))F_{L}(k-k^{\prime})$
$=\frac{4}{(2\pi)^{2d}}\int d^{d}q\Xi(q)F_{L}(q),$ (72)
where
$\Xi(q):=\int d^{d}k\eta(k)[1-\eta(q+k)].$ (73)
Similarly, we can bound the function $\Xi(q)$ with $s_{-}\|q\|_{2}$ where
$s_{-}=\frac{1}{2}A_{\text{FS}}$ (74)
with $A_{\text{FS}}$ defined as the area of the Fermi surface since the condition for $\text{Re}E_{k}=0$ and that for $\text{Im}E_{k}=0$ are the same here. Therefore, we obtain
$\int d^{d}qs_{-}\|q\|_{2}F_{L}(q)$ $\geq\frac{1}{\sqrt{d}}\int d^{d}qs_{-}\|q\|_{1}F_{L}(q)$
$=\sum_{i=1}^{d}\frac{1}{\sqrt{d}}\int d^{d}qs_{-}F_{L}(q)q_{i}$
[…]
$S_{A}>\frac{2\sqrt{d}}{(2\pi)^{d+1}}s_{-}L^{d-1}\log L:=c_{-}L^{d-1}\log L,$ (76)
where $c_{-}:=\frac{2\sqrt{d}}{(2\pi)^{d+1}}s_{-}$. The entanglement entropy shows at least a logarithmic scaling when $d=1$. For the multi-band cases, we have $S_{A}>\lfloor N/2\rfloor c_{-}\log L$, where $\lfloor x\rfloor$ is the floor function. Hence, the PT-unbroken phase with Yang-Lee zeros is always accompanied by at least a subarea-law scaling for the entanglement entropy. In the PT-broken phase, the entanglement entropy exhibits area law since it is gapped. Hence, the edges of the distribution of Yang-Lee zeros in one-dimensional cases always correspond to the entanglement transition. For the other cases where the Fermi surface is not the exceptional ring, the discussion is exactly the same as the Hermitian one [40]. Since Yang-Lee zeros can only take place in the gapless PT-broken phase, the edge of the Yang-Lee zeros can always correspond to the entanglement transitions.
Appendix C Derivation of Eqs. (16) and (17)
We here consider the one-dimensional (1D) XXZ model with the Hamiltonian
$H=-J\sum_{i=1}^{L}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}+\Delta S_{i}^{z}S_{i+1}^{z}),$ (77)
where $J>0$ and $\Delta$ is the anisotropy parameter. This model can be exactly solved by the Bethe ansatz. The Bethe ansatz equations of the XXZ model are given by [60]
$\left(\frac{\sin\frac{\phi}{2}(x_{j}+i)}{\sin\frac{\phi}{2}(x_{j}-i)}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l}+2i)}{\sin\frac{\phi}{2}(x_{j}-x_{l}-2i)}.$ (78)
Here we define $\cosh\phi:=\Delta$ and $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$, where $k_{j}$ is the quasimomentum. The index is given by $j=1,\cdots,M$ with $M$ being the number of the excited magnons. If we take the limit $\delta:=\Delta-1\rightarrow 0$, we can approximate $\phi$ as $\phi\simeq\sqrt{2\delta}$. When $\phi=0$ ($\Delta$ = 1), Eq. (78) can be reduced as
$\left(\frac{x_{j}+i}{x_{j}-i}\right)^{L}=\prod_{l\neq j}\left(\frac{x_{j}-x_{l}+2i}{x_{j}-x_{l}-2i}\right),$ (79)
which consists of the Bethe ansatz equations of the Heisenberg model. In this case the variable $x_{j}$ can be rewritten as $x_{j}=\cot(k_{j}/2)$. From Eq. (79), we can see that $x_{j}=\infty(k_{j}=0)$ is a set of solutions of the Bethe ansatz, which represents that all magnons have zero momentum or equal-weight superposition on each site of the system. These states are ground states of the Heisenberg model deduced from the energy:
$E_{M}=E_{0}+\sum_{j=1}^{M}J(1-\cos k_{j})=E_{0}+\sum_{j=1}^{M}\frac{2J}{x_{j}^{2}+1}.$ (80)
For the Heisenberg model with a system size $L$, there is $(L+1)$-degeneracy of the ground state, originating from the SU(2) symmetry of the model. These eigenstates can be expressed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where $S^{-}=\sum_{j}S_{j}^{-}$. However, when we have a finite $\phi$ for anisotropy, the degeneracy will be lifted. To be specific, we begin from the expansion of the Bethe ansatz equations (78) of the XXZ model. Since we only consider the distribution of Yang-Lee zeros around the phase transition point $\Delta=1$, we take the limit $\phi\rightarrow 0$ and expand the equation as
$\left(\frac{\sin\frac{\phi}{2}x_{j}+i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}{\sin\frac{\phi}{2}x_{j}-i\frac{\phi}{2}\cos\frac{\phi}{2}x_{j}}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l})+i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}{\sin\frac{\phi}{2}(x_{j}-x_{l})-i\phi\cos\frac{\phi}{2}(x_{j}-x_{l})}.$ (81)
Up to the first order of $\phi,$ this expansion can be rewritten as
$L\cot\frac{\phi}{2}x_{j}=2\sum_{l\neq j}\cot\frac{\phi}{2}(x_{j}-x_{l}),$ (82)
which is referred to as simplified Bethe ansatz equations (SBAEs). If we define $y_{j}=\frac{\phi}{2}x_{j}$, the SBAEs can be rewritten as
$L\cot y_{j}=2\sum_{l\neq j}\cot(y_{j}-y_{l}).$ (83)
We can see that this equation is purely an algebraic equation on $\cot y_{j}$, where the solutions only depend on the size $L$ and the magnon number $M$. Meanwhile, the energy of the $M$-magnon excitation is given by
$E=E_{0}+\sum_{j=1}^{M}\frac{J\sinh^{2}\phi}{\cosh\phi-\cos\phi x_{j}}$ $=$ $E_{0}+\sum_{j=1}^{M}\frac{2J\delta}{1+\delta-\cos\phi x_{j}}$ (84)
$=$ $E_{0}+\sum_{j=1}^{M}\frac{2J\delta(1+\cot^{2}y_{j})}{2+\delta\cot^{2}y_{j}}.$
In the expansion above, we first assume that $\cot y_{j}$ is finite so that we can expand the Bethe ansatz equation (78) up to the first order of $\phi$. Now we can also see this point from the expression (84). Divergence of the factor $\cot y_{j}$ will increase $2J$ to the energy, indicating that this solution does not originate from the original $(L+1)$-degeneracy of the Heisenberg spin chain. Therefore, we can exactly obtain all the deviated ground states from the SBAE. The finiteness of $\cot y_{j}$ of the deviated ground states can also be seen from the SBAE (82). We assume that there are divergences of $\cot y_{j}$ first. Then there must be a series of $\cot y_{j}$ which satisfy $\cot y_{j}=a_{j}N$, where $j=1,\cdots,m$, with $N$ being a divergent number and $a_{j}$ being a constant. Since some of the solutions $\cot y_{j}$ may not diverge, we always have $m\leqslant M\leqslant L$. Then the SBAE becomes
$La_{j}=2\sum_{l\neq j}\frac{a_{j}a_{l}}{a_{l}-a_{j}}.$ (85)
By summing up all the equalities, we have
$mL=2\frac{m(m-1)}{2}\Rightarrow L=m-1.$ (86)
However, this contradicts $m\leqslant L$. Therefore, all of the $\cot y_{j}$ should be convergent.
We then move to prove Eq. (16) in the main text. For convenience, we define a new variable $\zeta_{j}=\cot y_{j}$. Then the SBAE can be rewritten as
$L\zeta_{j}=2\sum_{l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}.$ (87)
Then we have
$\sum_{j}\zeta_{j}^{2}$ $=$ $\frac{2}{L}\sum_{j,l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}\zeta_{j}$ (88)
$=$ $-\frac{2}{L}\sum_{j,l\neq j}\frac{1+\zeta_{l}\zeta_{j}}{\zeta_{l}-\zeta_{j}}\zeta_{l},$
where we exchange the label $l,j$ in the second equality. Hence, by summation of the right-hand side of the two equalities, we obtain
$\sum_{j}\zeta_{j}^{2}$ $=$ $-\frac{1}{L}\sum_{j,l\neq j}(1+\zeta_{l}\zeta_{j})$ (89)
$=$ $-\frac{M(M-1)}{L}-\frac{1}{L}\sum_{j,l\neq j}\zeta_{l}\zeta_{j}$
$=$ $-\frac{M(M-1)}{L}-\frac{1}{L}\left(\sum_{j}\zeta_{j}\right)^{2}+\frac{1}{L}\sum_{j}\zeta_{j}^{2}$
$=$ $-\frac{M(M-1)}{L}+\frac{1}{L}\sum_{j}\zeta_{j}^{2},$
where we use $\sum_{j}\zeta_{j}=0$ from Eq. (87). Hence, the summation of $\zeta_{j}^{2}$ can be simplified as
$\sum_{j}\zeta_{j}^{2}=-\frac{M(M-1)}{L-1}.$
By revisiting the expression of energy (84), we have
$E=E_{0}+\sum_{j=1}^{M}J\delta(1+\cot^{2}y_{j})=E_{0}+J\delta\left(M+\sum_{j=1}^{M}\zeta_{j}^{2}\right)=E_{0}+J\delta\frac{M(L-M)}{L-1},$ (90)
which can indicate that the energy reaches the maximum for $M=\frac{L}{2}\left(\frac{L-1}{2}\right)$ when $L$ is even (odd). With the expression of energy (90), we show the expression of the partition function at absolute zero as
$Z=\lim_{T\rightarrow 0}\text{Tr}e^{-\beta H}=\lim_{T\rightarrow 0}\sum_{M=0}^{L}e^{-\beta E_{M}}=\lim_{T\rightarrow 0}e^{-\beta E_{0}}\sum_{M=0}^{L}z^{M(L-M)}.$ (91)
Hence, the condition of $Z=0$ is equivalent to
$\sum_{M=0}^{L}z^{M(L-M)}=0,$ (92)
which is exactly Eq. (17) or Eq. (10) in the main text. By solving this equation, we can obtain the distribution of Yang-Lee zeros, which is given by $\text{Re}\Delta=1$ on the complex plane of $\Delta$.
Finally, we discuss the Yang-Lee edge singularity associated with the Yang-Lee zeros. Around the phase transition point $\Delta=1$, we consider the susceptibility [60]
$\chi=\frac{2s_{z}}{h}=\frac{4\gamma}{J\pi(\pi-\gamma)\sin\gamma}[1+O(h^{2})+O(h^{4\gamma/(\pi-\gamma)})],$ (93)
where
$\cos\gamma=-\Delta.$ (94)
Around the ferromagnetic phase transition point, we have $\gamma=\pi$, which shows the divergence of the susceptibility. The scaling of $\gamma$ near $\Delta=1$ is $\gamma\approx\pi-i\sqrt{2\delta}$. Then we generalize $\delta$ into an imaginary number where Yang-Lee zeros are located. The scaling of the susceptibility is given by
$\chi(h=0)\simeq\frac{4\gamma}{J\pi\left(-i\sqrt{2\delta}\right)i\sqrt{2\delta}}=\frac{2}{J\delta}\propto\delta^{-1}.$ (95)
Furthermore, we calculate the correlation length. Due to the finite deviation $\delta$, the ground-state degeneracy is lifted and the system becomes gapped. Near the ferromagnetic phase transition point, the gap $\Delta E$ is given by
$\Delta E=\text{Re}[J\delta],$ (96)
which is determined by the Bethe ansatz. Hence, from the solution we can see the critical exponent $\varphi=1$, where we define $\Delta E=J\times\text{Re}[\delta]^{\varphi}$.